OpenStudy (anonymous):

3 over 12 divided by 5 over 6 as a mixed number

OpenStudy (anonymous):

(3/12)/(5/6)

OpenStudy (anonymous):

Idk

OpenStudy (anonymous):

\[\frac{3}{12} \div \frac{5}{6}=\frac{3}{12} \times \frac{6}{5} =\frac{3}{2} \times \frac{1}{5} =\frac{3}{10}\]